How do you set up the proportion for, what percent of 52 is 13
Found 3 solutions by rfer, stanbon, Edog1326:
Answer by rfer(16322)

13/54=0.24=24%

Answer by stanbon(75887)

set up the proportion for, what percent of 52 is 13
-----
x/100 = 13/52
----
x = 100(13/52) = 1300/52 = 25
----
Answer: 25%
=================
Cheers,
Stan H.
==========

Answer by Edog1326(1)

A fraction is set up like this:
part/whole
A percent fraction is set up like this:
part/100 (since a percent is out of 100)
Then set it up as a proportion. The proportion would be set up like this:
part/whole(fraction)=part/100 (percent)
So then the actual proportion would look like this...
13/52=x/100.
x is the variable or missing part of your proportion.